Megan is having her yard landscaped. She obtained an estimate from two landscaping companies. Company A gave an estimate of $240 for materials and equipment rental plus $55 per hour for labor. Company B gave an estimate of $320 for materials and equipment rental plus $45 per hour for labor. Determine how many hours of labor will be required for the two companies to cost the same.

Answer by ikleyn(52915)

.
    Company A    Company B


    240 + 55*h = 320 + 45*h


Making equation is so simple, that does not require explanations.


Solving this equation is very simple, too.  

You collect and combine the terms with "h" on the left side, keeping the constant terms on the right side


    55h - 45h = 320 - 240

       10h    = 80

         h    = 80/10 = 8 hours.    ANSWER

Solved.

--------------

Is everything clear to you now ?

Are you fully satisfied ?

Your satisfaction is important for us . . .